Salvatore Montalbano is a shareholder in the Enel Corporation. The current price of Enel’s stock is $36 per share, and there are 1 million shares outstanding. Salvatore owns 20,000 shares, or 2% of the stock, which she purchased one year ago for $35 per share.

Assume that Enel makes a surprise announcement that it plans to repurchase 100,000 shares of its own stock at a price of $40 per share. In response to this announcement, Enel’s stock price increases $2 per share, from $36 to $38, but this price is expected to fall back to $37 per share after the repurchase is completed.

Assume that Salvatore faces marginal personal tax rates of 15% on both dividend income and capital gains.

a) Calculate Salvatore’s (realized) after-tax return in dollar terms and percentage terms from his investment in Enel shares, assuming that he chooses to participate in the repurchase program and all of the shares she tenders are purchased at $40 per share.

b) How many shares will Salvatore be able to sell if all Enel shareholders tender their shares to the firm as part of this repurchase program and the company purchases shares on a pro rata basis?

c) What fraction of Enel’s total common equity will Salvatore own after the repurchase program is completed if she chooses not to tender his shares?

Please show your calculations clearly
Answer #1

Given information:

Salvatore owns 20,000 shares, or 2% of the stock, purchase price $35 per share.

current price of Enel’s stock is $36 per share

Enel plans to repurchase 100,000 shares at a price of $40 per share

a) Salvatore's after tax returns can be calculated as below

20,000 shares * (40-35)=$100,000

After tax profit would be $100,000 (1-15%) = $85,000

b) If Enel purchases shares at a pro rata basis, Salvatore will be able to sell only 2% of her 20,000 share

which is 20,000 * 2%= 400 shares

c) IF Salvatore choses not participate in the programe after the buy back she will be owning can be caluclated using the below formula

Salvatore's shares/(Current outstanding of Enel + Repurchase )

20,000/(1,000,000+100,000) = 1.82%